# DeepPAAC: A New Deep Galerkin Method for Principal-Agent Problems

### v1

@@ 4.1 A Solely Continuous Payment
The Principal’s value function $V^{P}$ satisfies the following HJB equation:
$\begin{cases}V^{P}_{t}+\sup_{\alpha,\beta,Z}\left\{(-C_{0}+{1\over 2}Z^{2}-{1\over 2}\beta^{2}-\alpha)V^{P}_{w}+{1\over 2}Z^{2}V^{P}_{ww}+(1-\beta)(\beta+Z)-\alpha\right\}=0,\\ V^{P}(T,w)=-w\end{cases}$ (4.28)
@@
To solve (4.28) numerically, we proceed to set up two different neural networks: one for approximating the value function $V^{P}$, and the other for jointly modeling the control processes $\alpha,\beta,Z$ (an alternative, discussed below is to train 3 separate networks for each of $\alpha,\beta,Z$).
The first order conditions for each control are:
${\cal F}^{\alpha}:=-V^{P}_{w}-1$
${\cal F}^{\beta}:=-\beta V^{P}_{w}+1-2\beta-Z$ (4.29)
${\cal F}^{Z}:=ZV^{P}_{w}+ZV^{P}_{ww}+1-\beta$
We additionally check that all the above f.o.c. are approximately satisfied, in other words all the above RHS ${\cal F}^{\alpha,\beta,Z}\simeq 0$ are close (in $L_{\infty}$ norm) to zero based on the numerically fitted neural networks for $V$ and for $\alpha,\beta,Z$, summarized as $L_{ctrl}$. For NN hyperparameters, we take training batches of $M=2000$, with $B=10$ steps per epoch. The polynomial learning rate decays from $10^{-3}$ to $10^{-4}$ with power $0.8$. The stopping rule is $\mathrm{Tol_{int}}=10^{-2}$ and $\mathrm{Tol_{ctrl}}=10^{-3}$.
Choosing $C_{0}=0$, we solve (4.28); numerical convergence to above tolerance is achieved in 300 training steps. Figure 4 shows the pointwise HJB residuals ${\bar{L}}_{int}(\theta_{n}^{V};t,x)$ as a function of $(t,x)$ at $n=300$, as well as the loss metrics $\|{\bar{L}}_{int}(\theta_{n}^{V};t_{n}^{\cdot},\bar{{\bf x}}_{n}^{\cdot})\|_{2}$ and $\|{\bar{L}}_{int}(\theta_{n}^{V};t_{n}^{\cdot},\bar{{\bf x}}_{n}^{\cdot})\|_{\infty}$ as a function of training steps $n$.
Figure 4: Right (a): Contour plot of the error metrics $\bar{L}_{int}(\theta_{n}^{V};t_{n}^{m},\bar{{\bf x}}_{n}^{\cdot})$ at $n=300$ training steps. Left (b): $L_{2}$-norm and $L_{\infty}$-norm for $\bar{L}_{int}(\theta_{n}^{V};t_{n}^{m},\bar{{\bf x}}_{n}^{\cdot})$ across training steps $n$.
@@ 4.2 DeepPAAC Variants
Table 1: Relative performance of different algorithms for Example of Section 4.1. We report results across 5 runs to validate algorithm stability. The second column shows median number of SGD steps to achieve convergence and the third column the respective range across the 5 runs. Running time is in seconds. In the first row, one run never achieved the stopping threshold over 10,000 steps.
@@
Table 1 shows that learning the two controls using one shared NN performs better than utilizing two separate NNs in our DeepPAAC algorithm. Moreover, our algorithm converges faster than the DGM architecture. Figure 5 plots the $L_{\infty}$-norm for $\bar{L}_{int}(\theta^{V}_{n};{\bar{t}}^{\cdot}_{n},{\bar{{\bf x}}}^{\cdot}_{n})$, and $\bar{L}_{ctrl}(\theta^{V}_{n},{\bar{t}}^{\cdot}_{n},{\bar{{\bf x}}}^{\cdot}_{n})$ of using DeepPAAC with a single shared NN for learning $\bf u$}.
(a) $\|\bar{L}_{int}(\theta^{V}_{n};{\bar{t}}^{\cdot}_{n},{\bar{{\bf x}}}^{\cdot}_{n})\|_{\infty}$ across 5 runs
(b) $\|\bar{L}_{ctrl}(\theta^{V}_{n};{\bar{t}}^{\cdot}_{n},{\bar{{\bf x}}}^{\cdot}_{n})\|_{\infty}$ across 5 runs
Figure 5: Validating convergence of the DeepPAAC loss functions. We show the $L_{\infty}$-norm for the value function HJB residual $\bar{L}_{int}(\theta^{V}_{n};{\bar{t}}^{\cdot}_{n},{\bar{{\bf x}}}^{\cdot}_{n})$ (left panel (a)) and the control loss $\bar{L}_{ctrl}(\theta^{V}_{n}{\bar{t}}^{\cdot}_{n},{\bar{{\bf x}}}^{\cdot}_{n})$ (right panel (b)) as a function of step $n$, for the case study from Section 4.1. Five runs of the DeepPAAC scheme with a single control network.
@@
  • Residual-based adaptive sampling: one strategy to speed up convergence is to preferentially train on samples where the loss functions are largest [wu2023comprehensive]. To test this, at each epoch we generate $4M$ candidate points in $(0,T)\times{\cal X}$. After sorting them by the magnitude of the PDE residual magnitude $L_{1}(t,{\bf x})$ we keep the top 20% of the points (the “high-error region”) which yield $0.8M$ samples, as well as retain a random sub-sample of $0.2M$ from the remaining “low-error” candidates.
@@
$M$ effectively acts as our batch size; we find that increasing the number of sampling points generally accelerates convergence, with only a marginal increase in the running time per training step. This happens because modern libraries distribute the loss function evaluation across $M$ training locations across the GPU/CPU memory. As the number of samples increases, the workload raises hardware utilization but there is minimal wall-clock impact until the memory resources are exhausted. At the same time, a large $M$ leads to a more accurate (i.e., less noisy) computation of the gradients and hence easier navigation of the loss landscape towards its minimum. As such, we recommend taking $M$ as big as possible given the computational capacity and memory resources of the hardware employed during training.
@@ 4.3 PA problem with Constrained Contracts
Under constraints (4.30), the Principal’s value function $V^{P}$ satisfies the HJB equation:
$\begin{cases}V^{P}_{t}+\sup\limits_{\begin{subarray}{c}\alpha,\ \underline{\beta}\leq\beta\leq\bar{\beta},\\ \underline{u}\leq\beta+Z\leq\overline{u}\end{subarray}}\left\{(-C_{0}+{1\over 2}Z^{2}-{1\over 2}\beta^{2}-\alpha)V^{P}_{w}+{1\over 2}Z^{2}V^{P}_{ww}+(1-\beta)(\beta+Z)-\alpha\right\}=0\cr V^{P}(T,w)=-w,\end{cases}$ (4.31)
@@
$L_{u}(\theta_{n}^{u})=$ $-\sum_{m=1}^{M}\Bigl[{\cal L}^{{\bf u}(t_{n}^{m},{\bf x}_{n}^{m};\theta_{n}^{u})}v(t_{n}^{m},{\bf x}_{n}^{m};\theta_{n}^{V})+v(t_{n}^{m},{\bf x}_{n}^{m},{\bf u}(t_{n}^{m},{\bf x}_{n}^{m};\theta_{n}^{u}))$
@@
In our example above, since the constraints are on $\beta$ and $Z$, the first order condition becomes:
${\cal F}^{\alpha}:=-V^{P}_{w}-1.$ (4.35)
To solve (4.31) numerically, we use the architecture design as in Section 4.1. The stopping rule is $\mathrm{Tol_{int}}=10^{-2},\mathrm{Tol_{ctrl}}=10^{-3}$ and $\overline{P}_{\infty}=0$ with $M=2000$ training points and $B=10$ steps per epoch. Learning rate schedule is with polynomial decay from $10^{-3}$ to $10^{-4}$ with power $0.8$. In the following three numerical examples, we illustrate several different constraint combinations. Throughout we choose $C_{0}=0$.
@@
In top row of Figure 7, we plot the numerically obtained optimal controls $\hat{\beta}(t,x)$, as well as their sum $\hat{\beta}(t,x)+\hat{z}(t,x)$ (so one can read off $\hat{z}$ as well). We observe that the constraint on $\beta+Z$ is binding, $\hat{\beta}(t,x)+\hat{z}(t,x)\simeq\overline{u}=0.5$ as expected. But the constraint of $\hat{\beta}(t,x)$ is not binding. Panel (c) shows the $L_{1}$ and $L_{\infty}$ norms of the penalty function on the validation set:
$\overline{P}_{1}:=\frac{1}{M}\sum_{m=1}^{M_{V}}\overline{P}(t_{n}^{m},{\bf x}_{n}^{m},u(t_{n}^{m},{\bf x}_{n}^{m};\theta_{n}^{u})),\qquad\overline{P}_{\infty}:=\max_{m}\overline{P}(t_{n}^{m},{\bf x}_{n}^{m},u(t_{n}^{m},{\bf x}_{n}^{m};\theta_{n}^{u}))$
as a function of steps $n$ until convergence.
(a) $\hat{\beta}(t,w)$
(b) $\hat{\beta}(t,w)+\hat{z}(t,w)$
(c) Loss function
(d) $\hat{\beta}(t,w)$
(e) $\hat{\beta}(t,w)+\hat{z}(t,w)$
(f) Loss function
(g) $\hat{\beta}(t,w)$
(h) $\hat{\beta}(t,w)+\hat{z}(t,w)$
(i) Loss function
Figure 7: Validating the solution of the constrained PA case study from Section 4.3 across three different constraint settings. Top row: $0\leq\beta\leq 0.1,0\leq\beta+Z\leq 0.5$. Middle row: $0\leq\beta\leq 0.2$. Bottom row: $\beta\leq 0.5,\beta+Z\leq 1.2$. The right-most panels (c)(f)(i)(c)-(f)-(i) trace the $L_{2}$-norms and $L_{\infty}$-norms for $\bar{L}_{int}(\theta_{n}^{V};t_{n}^{m},\bar{{\bf x}}_{n}^{m})$ as well as for the penalty functions $\overline{P}_{1}$, $\overline{P}_{\infty}$. Numerical convergence is achieved in $350/380/430$ training steps for cases 1-3 respectively.
@@
The middle row of Figure 7 shows that the constraint is no longer binding in the interior, and $\hat{\beta}(t,x)$ can be as low as 0.16 for large $t$.
@@
The bottom row of Figure 7 shows that (as expected) the constraints on $\beta$ and $\beta+Z$ do not bind, and the sum $\hat{\beta}(t,x)+\hat{z}(t,x)$ remains 1 everywhere.
@@ 4.4 Two-dimensional Problem with Scalar Control and Explicit Solution
$U^{A}(t,\alpha,\beta,a,c)=-{1\over 2}a^{2},\quad U^{P}(t,\alpha,\beta,a,D)=0.$
For the terminal lump-sum payment, there is a risk-averse Agent utility function $\Phi^{A}(T,\xi,m)=\log\xi$ and risk-neutral Principal’s utility function $\Phi^{P}(T,\xi,X,m)=X-\xi$.
The Agent’s optimal wealth process $W_{t}$ satisfies
@@
$V^{P}(t,x,w)=x-e^{w}+{1\over 4}e^{-w}x^{2}\big(e^{\sigma^{2}(T-t)}-1\big),\qquad Z^{*}={1\over 2}e^{-w}.$
@@
For the NN training, we use $M=2000$ training points per epoch and $B=10$ SGD steps. The learning rate schedule is of polynomial type with power $0.8$, $\alpha_{0}=10^{-3}$ decaying to $\alpha_{10^{5}}=10^{-5}$. With the stopping rule $\mathrm{Tol_{int}}=\mathrm{Tol_{ctrl}}=10^{-3}$, convergence is achieved in $51570$ training iterations, see Figure 8.
(a) absolute error for value function at $t=0$
(b) absolute error for optimal control at $t=0$
Figure 8: Heatmaps of the absolute errors in the value function $V(0,x,w)$ and optimal control $a(0,x,w)$ at $t=0$ relative to the ground truth for the example of Section $4.$

### v2

@@ 4.1 A Solely Continuous Payment
The Principal’s value function $V^{P}(t,w)$ satisfies the following HJB equation:
$\begin{cases}V^{P}_{t}+\sup\limits_{\alpha,\beta,Z}\left\{(-C_{0}+{1\over 2}Z^{2}-{1\over 2}\beta^{2}-\alpha)V^{P}_{w}+{1\over 2}Z^{2}V^{P}_{ww}+(1-\beta)(\beta+Z)-\alpha\right\}=0,\\ V^{P}(T,w)=-w\end{cases}$ (4.29)
@@
To solve (4.29) numerically, we proceed to set up two different neural networks: one for approximating the value function $V^{P},$ and the other for jointly modeling the control processes $\alpha,\beta,Z$ (an alternative, discussed below, is to train 3 separate networks for each of $\alpha,\beta,Z$).
For this example, we have the vector convergence loss criterion
$L_{ctrl}=\Big[-V^{P}_{w}-1,\;-\beta V^{P}_{w}+1-2\beta-Z,\;ZV^{P}_{w}+ZV^{P}_{ww}+1-\beta\Big].$ (4.30)
For NN hyperparameters, we take training batches of $M=2000$, with $B=10$ steps per epoch. The polynomial learning rate decays from $10^{-3}$ to $10^{-4}$ with power $0.8$. The stopping rule is $\mathrm{Tol_{int}}=10^{-2}$ and $\mathrm{Tol_{ctrl}}=10^{-3}$, since the quantity $\|\bar{L}_{ctrl}(\theta^{V}_{n},\theta^{u}_{n};{\bar{t}}^{\cdot},{\bar{w}}^{\cdot})\|_{\infty}$ typically decays faster than the loss $\|\bar{L}_{int}(\theta^{V}_{n};{\bar{t}}^{\cdot},\bar{w}^{\cdot})\|_{\infty}$ in our experiments.
Choosing $C_{0}=0$, we solve (4.29); numerical convergence to above tolerance is achieved in 300 training steps. Figure 4 shows the pointwise HJB residuals ${\bar{L}}_{int}(\theta_{n}^{V};t,w)$ as a function of $(t,w)$ at $n=300$, as well as the $L_{2}$ and $L_{\infty}$-norm of the validation loss metric ${\bar{L}}_{int}(\theta_{n}^{V};\bar{t}^{\cdot},\bar{w}^{\cdot})$ as a function of training steps $n$.
Figure 4: Left (a): Contour plot of the value function error metrics $\bar{L}_{int}(\theta_{n}^{V};t,w)$ at $n=300$ training steps. Right (b): $L_{2}$-norm and $L_{\infty}$-norm for $\bar{L}_{int}(\theta_{n}^{V};\bar{t}^{\cdot},\bar{w}^{\cdot})$ across $n$.
@@ 4.2 DeepPAAC Variants
Table 1: Relative performance of different algorithms for case study of Section 4.1. We report results across 5 runs to validate algorithm stability. The second column shows median number of SGD steps to achieve convergence and the third column the respective range across the 5 runs. Running time is in seconds. In the first row, one run never achieved the stopping threshold over 10,000 steps.
@@
Table 1 shows that learning the three controls ${\bf u}=(\alpha,\beta,Z)$ using a common shared NN performs better than utilizing three separate NNs in our DeepPAAC algorithm. Moreover, our algorithm converges faster than the DGM architecture. Figure 5 plots the $L_{\infty}$-norm for $\bar{L}_{int}(\theta^{V}_{n};{\bar{t}}^{\cdot},\bar{w}^{\cdot})$, and $\bar{L}_{ctrl}(\theta^{V}_{n},\theta^{u}_{n};{\bar{t}}^{\cdot},\bar{w}^{\cdot})$ of using DeepPAAC with a single shared NN for learning $\bf u$}.
(a) $\|\bar{L}_{int}(\theta^{V}_{n};{\bar{t}}^{\cdot},\bar{w}^{\cdot})\|_{\infty}$ across 5 runs
(b) $\|\bar{L}_{ctrl}(\theta^{V}_{n},\theta^{u}_{n};{\bar{t}}^{\cdot},{\bar{w}}^{\cdot})\|_{\infty}$ across 5 runs
Figure 5: Validating convergence of the DeepPAAC loss functions. We show the $L_{\infty}$-norm for the value function HJB residual $\bar{L}_{int}(\theta^{V}_{n};{\bar{t}}^{\cdot},\bar{w}^{\cdot})$ (left panel (a)) and the control loss criterion $\bar{L}_{ctrl}(\theta^{V}_{n},\theta^{u}_{n};{\bar{t}}^{\cdot},\bar{w}^{\cdot})$ (right panel (b)) as a function of step $n$, for the case study from Section 4.1. Five runs of the DeepPAAC scheme with a single control $N$.
@@
  • Residual-based adaptive sampling (RAD): one strategy to speed up convergence is to preferentially train on samples where the loss functions are largest [wu2023comprehensive]. To test this, at each epoch we generate $4M$ candidate points in $(0,T)\times{\cal X}$. After sorting them by the magnitude of the PDE residual magnitude $L_{1}(t,w)$ we keep the top 20% of the points (the “high-error region”) which yield $0.8M$ samples, as well as retain a random sub-sample of $0.2M$ from the remaining “low-error” candidates.
@@
The parameter $M$ effectively acts as our batch size; we find that increasing the number of sampling points generally accelerates convergence, with only a marginal increase in the running time per training step. This happens because modern libraries distribute the loss function evaluation across $M$ training locations across the GPU/CPU memory. As the number of samples increases, the workload raises hardware utilization but there is minimal wall-clock impact until the memory resources are exhausted. At the same time, a large $M$ leads to a more accurate (i.e., less noisy) computation of the gradients and hence easier navigation of the loss landscape towards its minimum. As such, we recommend taking $M$ as big as possible given the computational capacity and memory resources of the hardware employed during training.
@@ 4.3 PA problem with Constrained Contracts
Under constraints (4.31), the Principal’s value function $V^{P}(t,w)$ satisfies the HJB equation:
$\hskip 20.0pt\begin{cases}V^{P}_{t}+\sup\limits_{\begin{subarray}{c}\alpha,\ \underline{\beta}\leq\beta\leq\bar{\beta},\\ \underline{u}\leq\beta+Z\leq\overline{u}\end{subarray}}\left\{(-C_{0}+{1\over 2}Z^{2}-{1\over 2}\beta^{2}-\alpha)V^{P}_{w}+{1\over 2}Z^{2}V^{P}_{ww}+(1-\beta)(\beta+Z)-\alpha\right\}=0\cr V^{P}(T,w)=-w,\end{cases}$ (4.32)
@@
$L_{u}(\theta_{n}^{u})=$ $-\sum_{m=1}^{M_{V}}\Bigl[{\cal L}^{{\bf u}(t_{n}^{m},{\bf x}_{n}^{m};\theta_{n}^{u})}v(t_{n}^{m},{\bf x}_{n}^{m};\theta_{n}^{V})+v(t_{n}^{m},{\bf x}_{n}^{m},{\bf u}(t_{n}^{m},{\bf x}_{n}^{m};\theta_{n}^{u}))$
@@
In our example above, since the constraints are on $\beta$ and $Z$, we have the convergence loss criterion reducing to
$L_{ctrl}=\Big[-V^{P}_{w}-1\Big].$ (4.36)
(a) ${\beta}(t,w)$
(b) ${\beta}(t,w)+{Z}(t,w)$
(c) Loss function
(d) ${\beta}(t,w)$
(e) ${\beta}(t,w)+{Z}(t,w)$
(f) Loss function
(g) ${\beta}(t,w)$
(h) ${\beta}(t,w)+{Z}(t,w)$
(i) Loss function
Figure 7: Validating the solution of the constrained PA case study from Section 4.3 across three different constraint settings. Top row: $0\leq\beta\leq 0.1,0\leq\beta+Z\leq 0.5$. Middle row: $0\leq\beta\leq 0.2$. Bottom row: $\beta\leq 0.5,\beta+Z\leq 1.2$. The right-most panels (c)(f)(i)(c)-(f)-(i) trace the $L_{2}$-norms and $L_{\infty}$-norms of $\bar{L}_{int}(\theta_{n}^{V};\bar{t}^{\cdot},\bar{w}^{\cdot})$ as well as the penalty functions $\overline{P}_{1}$, $\overline{P}_{\infty}$. Numerical convergence is achieved in $350/380/430$ training steps for cases 1-3 respectively.
To solve (4.32) numerically, we use the architecture design as in Section 4.1. The stopping rule is $\mathrm{Tol_{int}}=10^{-2},\mathrm{Tol_{ctrl}}=10^{-3}$ and $\overline{P}_{\infty}=0$ where the latter is the norm of the penalty function on the validation set $\{\bar{t}^{m},\bar{w}^{m}\}_{m=1}^{M_{V}}$:
$\overline{P}_{1}:=\frac{1}{M_{V}}\sum_{m=1}^{M_{V}}{P}(\bar{t}^{m},\bar{w}^{m},u(\bar{t}^{m},\bar{w}^{m};\theta_{n}^{u})),\qquad\overline{P}_{\infty}:=\max_{m\leq M_{V}}{P}(\bar{t}^{m},\bar{w}^{m},u(\bar{t}^{m},\bar{w}^{m};\theta_{n}^{u})).$
We use $M=2000$ training points and $B=10$ steps per epoch. Learning rate schedule is with polynomial decay from $10^{-3}$ to $10^{-4}$ with power $0.8$. In the following three numerical examples, we illustrate several different constraint combinations. Throughout we choose $C_{0}=0$.
@@
In top row of Figure 7, we plot the numerically obtained optimal control ${\beta}(t,w)$, as well as the sum ${\beta}(t,w)+{Z}(t,w)$ (so one can read off Z{Z} as well). We observe that the constraint on $\beta+Z$ is binding, ${\beta}(t,w)+{Z}(t,w)\simeq\overline{u}=0.5$ as expected. But the constraint of ${\beta}(t,w)$ is not binding. Panel (c) shows the $L_{1}$ and $L_{\infty}$ norms of the penalty function $\overline{P}_{1},\overline{P}_{\infty}$ as a function of steps $n$ until convergence.
@@
The middle row of Figure 7 shows that the constraint is no longer binding in the interior, and ${\beta}(t,w)$ can be as low as 0.16 for large $t$.
@@
The bottom row of Figure 7 shows that (as expected) the constraints on $\beta$ and $\beta+Z$ do not bind, and the sum ${\beta}(t,w)+Z(t,w)$ remains 1 everywhere.
@@ 4.4 Two-dimensional Problem with Scalar Control and Explicit Solution
$U^{A}(t,\alpha,\beta,a,c)=-{1\over 2}a^{2},\qquad U^{P}(t,\alpha,\beta,a,D)=0.$
For the terminal lump-sum payment, there is a risk-averse Agent utility function
$\Phi^{A}(T,\xi,m^{A})=\log\xi$ and risk-neutral Principal’s utility function $\Phi^{P}(T,\xi,X,m^{P})=X-\xi$.
The Agent’s optimal wealth process $\{W_{t}\}$ satisfies
@@
$V^{P}(t,x,w)=x-e^{w}+{1\over 4}e^{-w}x^{2}\big(e^{\sigma^{2}(T-t)}-1\big),\qquad Z^{*}(t,x,w)={1\over 2}e^{-w}.$
For this example, the convergence loss criterion is the gradient in $Z$ of the Hamiltonian:
$L_{ctrl}=\Big[(V^{P}_{w}+V^{P}_{ww})Z+V^{P}_{x}+V^{P}_{xw}\Big].$
@@
For the NN training, we use $M=2000$ training points per epoch and $B=10$ SGD steps. The learning rate schedule is of polynomial type with power $0.8$, $\alpha_{0}=10^{-3}$ decaying to $\alpha_{10^{5}}=10^{-5}$. With the stopping rule $\mathrm{Tol_{int}}=\mathrm{Tol_{ctrl}}=10^{-3}$, convergence is achieved in $51570$ training iterations for our DeepPAAC algorithm, whereas the DGM algorithm converges only after $94380$ training iterations, see Figure 8. Moreover, even after convergence the resulting absolute error compared to the exact ground truth solution are much worse: e.g. for $(x,w)$ in the neighborhood of $(1,0)$ (bottom right corner in Figure 8) the principal’s value function of DGM-PIA has error of >0.06>0.06 compared to just $0.0025$ for DeepPAAC.
(a) absolute error of DeepPAAC value function
(b) absolute error of DeepPAAC optimal control
(c) absolute error of DGM-PIA value function
(d) absolute error of DGM-PIA optimal control
Figure 8: Heatmaps of the absolute errors in the value function $V(0,x,w)$ and optimal control $a(0,x,w)$ at $t=0$ relative to the ground truth for the case study of Section $4.$ Top row: DeepPAAC algorithm. Bottom row: DGM-PIA from [al2019applications, al2022extensions].